# Spatially Regularized Super-Resolved Constrained Spherical Deconvolution (SR2-CSD) of Diffusion MRI Data

### v1

@@ 2.1 Spatially Regularized Super-Resolved CSD (SR2-CSD)
where $\boldsymbol{f}_{0}$ denotes a prior estimate of the FOD SH coefficients, computed by incorporating spatial information, and the non-negative parameter $\rho$ controls the weighting between the data-fitting and regularization terms. Notably, Eq. (8) can be reformulated as a QP problem similar to Eq. (7), but with parameters $\boldsymbol{K}$=$-\boldsymbol{P}$, $\boldsymbol{q}=\boldsymbol{0}$, $\boldsymbol{Q}$=$\boldsymbol{A}^{T}\boldsymbol{A}+\rho\boldsymbol{I}$, and $\boldsymbol{c}^{T}=-\boldsymbol{s}^{T}\boldsymbol{A}-\rho\boldsymbol{f}_{0}^{T}$.
@@
In practice, the TV filter [41] is applied independently to each 3D image formed by the FOD SH coefficients of a given degree and order. The strength of the filtering, defined by $\lambda=K\sigma$, varies depending on the noise standard deviation $\sigma$ affecting each SH 3D map. This noise standard deviation is estimated using a robust wavelet-based estimator [42], implemented in the estimate_sigma function available in the scikit-image Python package [43]. In the results section, we demonstrate the impact of the solution as a function of $K$, which is kept constant for all SH coefficients. However, since the optimal $K$ should depend on the complexity of the underlying anatomy of the image, it is generally not feasible to define a constant $K$ value that is optimal across different datasets and phantoms. To address this, we employ the J-invariance principle [44, 45] (see Section 2.2) to calibrate the parameter $K$ automatically. In our implementation, TV denoising is implemented using the Chambolle method [46], available in the scikit-image package. The smoothed FOD SH coefficients are then projected onto the constrained non-negative set, ensuring all the SH coefficients produce non-negative FOD amplitudes. This projection involves calculating the FOD amplitudes, setting negative values to zero, and recomputing the SH coefficients from the thresholded FODs. The resulting prior $\boldsymbol{f}_{0}$ is then used to solve the QP problem in Eq. (8). The regularization parameter is chosen as $\rho=1$. The pipeline is implemented in the DIPY software [38].
@@ 2.2 J-Invariance for Automatic TV Denoiser Calibration
The automatic TV denoiser calibration uses a self-supervised approach based on the Noise2Self algorithm [44]. A denoising function is considered J-invariant if the value of each pixel in the denoised output does not depend on the value of the same pixel in the noisy input image. If a function is not J-invariant, a masking procedure can be applied to make it J-invariant [44]. For a J-invariant denoiser, the self-supervised loss is equivalent to the ground-truth loss, up to a constant. This means that the difference between the denoised image and the noisy image can be used as a proxy for the difference between the denoised image and the original clean image. Therefore, we can calibrate a J-invariant denoiser using self-supervised loss, even without clean reference images. In our experiments, we searched for the optimal TV filtering strength $K$ within the range [0, 5].
@@ 3.5 Preprocessing
The Test-Retest data has already been preprocessed for motion, and measurements of the B0 magnetic field were performed to mitigate the effects of EPI spatial distortions, as detailed in the Stanford Digital Repository available at https://purl.stanford.edu/rt034xr8593. Despite this preprocessing, the two images collected per subject in this dataset could not be optimally spatially registered using a linear registration method due to residual spatial distortions affecting each image differently. Advanced tools like topup [54, 55] and eddy [56] in FSL could not be applied to further attenuate these distortions because no $b=0$ images were acquired with opposing polarities of the phase-encode blips. Instead, the estimated FODs for each subject from the two scans were spatially normalized using a non-linear registration algorithm, as described by [57], to achieve better spatial agreement.
@@
The response functions for SSST-CSD and MSMT-CSD experiments were estimated using DIPY. For SSST-CSD, the single-fiber response function was estimated using the response_from_mask_ssst function from the top 200 voxels with the highest fractional anisotropy. For MSMT-CSD, the response functions for WM, GM, and CSF tissue types were estimated using the mask_for_response_msmt function with default parameters as described in the DIPY documentation https://docs.dipy.org/stable/examples_built/reconstruction/reconst_mcsd.html (i.e., wm_fa_thr=0.7, gm_fa_thr=0.3, csf_fa_thr=0.15, gm_md_thr=0.001, csf_md_thr=0.0032).
@@ 4.1 Phantom
Subject MSE ACC AE PNE Subject MSE ACC AE PNE
Subject 1 Subject 4
CSD 0.265 0.867 19.30 0.188 CSD 0.266 0.871 19.01 0.191
Super-CSD 0.342 0.828 20.68 0.203 Super-CSD 0.337 0.835 20.10 0.208
SR2-CSD 0.180 0.910 17.53 0.184 SR2-CSD 0.173 0.918 16.70 0.184
Subject 2 Subject 5
CSD 0.277 0.861 20.07 0.190 CSD 0.234 0.883 19.75 0.209
Super-CSD 0.362 0.819 21.36 0.197 Super-CSD 0.316 0.842 20.99 0.217
SR2-CSD 0.179 0.912 18.09 0.188 SR2-CSD 0.179 0.910 17.86 0.206
Subject 3 Subject 6
CSD 0.295 0.852 19.65 0.182 CSD 0.325 0.837 20.74 0.194
Super-CSD 0.364 0.817 20.69 0.202 Super-CSD 0.407 0.796 21.74 0.205
SR2-CSD 0.193 0.903 17.33 0.175 SR2-CSD 0.224 0.888 18.83 0.203
@@ 5 Discussion and Conclusion
Future studies should explore alternative denoising methods for obtaining the spatial prior needed in SR2-CSD, including machine learning techniques like [59]. Additionally, our two-step analysis—computing the prior and re-estimating the FOD using the prior—could evolve into a more generalized iterative algorithm that minimizes a global multi-voxel cost function. This function would penalize the mean squared error between the measured and predicted data by the deconvolution model and the noise and lack of spatial continuity of the estimated FODs, potentially using a projected gradient descent method. However, we defer this implementation to future work, as incorporating the non-negativity constraint into this variant requires further investigation.
Future research should also evaluate whether the improved FOD estimates from SR2-CSD positively impact the connectivity estimation obtained by tractography methods [60], particularly regarding false positives and false negatives [61], as well as tracking reproducibility and subject-specificity [62]. The proposed method will be made freely available to the community in the DIPY software library [38].

### v2

@@ 2.1 Spatially Regularized Super-Resolved CSD (SR2-CSD)
where $\boldsymbol{f}_{0}$ denotes a prior estimate of the FOD SH coefficients, computed by incorporating spatial information, and the non-negative parameter $\rho$ controls the weighting between the data-fitting and regularization terms. Notably, Eq. (8) can be reformulated as a QP problem similar to Eq. (7), but with parameters $\boldsymbol{K}$=$-\boldsymbol{P}$, $\boldsymbol{q}=\boldsymbol{0}$, $\boldsymbol{Q}$=$\boldsymbol{A}^{T}\boldsymbol{A}+\rho\boldsymbol{I}$, and $\boldsymbol{c}^{T}=-\boldsymbol{s}^{T}\boldsymbol{A}-\rho\boldsymbol{f}_{0}^{T}$. Through empirical analysis, we determined that fixing the regularization parameter $\rho=1$ yields optimal performance across a range of datasets and noise levels.
@@
In practice, the TV filter [41] is applied independently to each 3D image formed by the FOD SH coefficients of a given degree and order. The strength of the filtering, defined by $\lambda=K\sigma$, varies depending on the noise standard deviation $\sigma$ affecting each SH 3D map. This noise standard deviation is estimated using a robust wavelet-based estimator [42], implemented in the estimate_sigma function available in the scikit-image Python package [43]. In the results section, we demonstrate the impact of the solution as a function of $K$, which is kept constant for all SH coefficients. However, since the optimal $K$ should depend on the complexity of the underlying anatomy of the image, it is generally not feasible to define a constant $K$ value that is optimal across different datasets and phantoms. To address this, we employ the J-invariance principle [44, 45] (see Section 2.2) to calibrate the parameter $K$ automatically. In our implementation, TV denoising is implemented using the Chambolle method [46], available in the scikit-image package. The smoothed FOD SH coefficients are then projected onto the constrained non-negative set, ensuring all the SH coefficients produce non-negative FOD amplitudes. This projection involves calculating the FOD amplitudes, setting negative values to zero, and recomputing the SH coefficients from the thresholded FODs. The resulting prior $\boldsymbol{f}_{0}$ is then used to solve the QP problem in Eq. (8). The pipeline is implemented in the DIPY software [38].
@@ 2.2 J-Invariance for Automatic TV Denoiser Calibration
The automatic TV denoiser calibration uses a self-supervised approach based on the Noise2Self algorithm [44]. A denoising function is considered J-invariant if the value of each pixel in the denoised output does not depend on the value of the same pixel in the noisy input image. If a function is not J-invariant, a masking procedure can be applied to make it J-invariant. For a J-invariant denoiser, the self-supervised loss is equivalent to the ground-truth loss, up to a constant [44]. This means that the difference between the denoised image and the noisy image can be used as a proxy for the difference between the denoised image and the original clean image. Therefore, we can calibrate a J-invariant denoiser using self-supervised loss, even without clean reference images. In our experiments, we searched for the optimal TV filtering strength $K$ within the range [0, 5].
@@ 3.5 Preprocessing
The Sherbrooke dataset was preprocessed for motion and eddy-current correction. The Test-Retest data has already been preprocessed for motion, and measurements of the B0 magnetic field were performed to mitigate the effects of EPI spatial distortions, as detailed in the Stanford Digital Repository available at https://purl.stanford.edu/rt034xr8593. Despite this preprocessing, the two images collected per subject in this dataset could not be optimally spatially registered using a linear registration method due to residual spatial distortions affecting each image differently. Advanced tools like topup [54, 55] and eddy [56] in FSL could not be applied to further attenuate these distortions because no $b=0$ images were acquired with opposing polarities of the phase-encode blips. Instead, the estimated FODs for each subject from the two scans were spatially normalized using a non-linear registration algorithm, as described by [57], to achieve better spatial agreement.
@@
The response functions for SSST-CSD and MSMT-CSD experiments were estimated using DIPY. For SSST-CSD, the single-fiber response function was estimated using the response_from_mask_ssst function from the top 200 voxels with the highest fractional anisotropy. For MSMT-CSD, the response functions for WM, GM, and CSF tissue types were estimated using the mask_for_response_msmt function with default parameters as described in the DIPY documentation https://docs.dipy.org/stable/examples_built/reconstruction/reconst_mcsd.html (i.e., wm_fa_thr=0.7, gm_fa_thr=0.3,
csf_fa_thr=0.15, gm_md_thr=0.001, csf_md_thr=0.0032).
@@ 4.3 Test-Retest dMRI Data
Subject MSE ACC AE PNE
Subject 1
CSD 0.265 0.867 19.30 0.188
Super-CSD 0.342 0.828 20.68 0.203
SR2-CSD 0.180 0.910 17.53 0.184
Subject 2
CSD 0.277 0.861 20.07 0.190
Super-CSD 0.362 0.819 21.36 0.197
SR2-CSD 0.179 0.912 18.09 0.188
Subject 3
CSD 0.295 0.852 19.65 0.182
Super-CSD 0.364 0.817 20.69 0.202
SR2-CSD 0.193 0.903 17.33 0.175
Subject MSE ACC AE PNE
Subject 4
CSD 0.266 0.871 19.01 0.191
Super-CSD 0.337 0.835 20.10 0.208
SR2-CSD 0.173 0.918 16.70 0.184
Subject 5
CSD 0.234 0.883 19.75 0.209
Super-CSD 0.316 0.842 20.99 0.217
SR2-CSD 0.179 0.910 17.86 0.206
Subject 6
CSD 0.325 0.837 20.74 0.194
Super-CSD 0.407 0.796 21.74 0.205
SR2-CSD 0.224 0.888 18.83 0.203
@@ 5 Discussion and Conclusion
Future studies should explore alternative denoising methods for obtaining the spatial prior needed in SR2-CSD, including machine learning techniques like [59]. Additionally, our two-step analysis—computing the prior and re-estimating the FOD using the prior—could evolve into a more generalized iterative algorithm that minimizes a global multi-voxel cost function. This function would penalize the mean squared error between the measured and predicted data by the deconvolution model and the noise and lack of spatial continuity of the estimated FODs, potentially using a projected gradient descent method. However, we defer this implementation to future work, as incorporating the non-negativity constraint into this variant requires further investigation. Moreover, the proposed method should be validated in real-world clinical conditions and populations with brain tissue abnormalities.
In this study, we used $l_{max}=12$ as it provides a good tradeoff between angular resolution and the number of estimated parameters. Notably, at least $l_{max}=12$ is required to operate in the super-resolution regime with the datasets used in this study. With this choice, we fit twice as many parameters compared to the conventional methods with $l_{max}=8$. Subsequent works could further investigate the optimal $l_{max}$ as a function of the number of diffusion gradient directions and SNRs. Nevertheless, our preliminary results showed similar performance with higher $l_{max}$ values (results not shown).
Further research should also evaluate whether the improved FOD estimates from SR2-CSD positively impact the connectivity estimation obtained by tractography methods [60], particularly regarding false positives and false negatives [61], as well as tracking reproducibility and subject-specificity [62]. The proposed method will be made freely available to the community in the DIPY software library [38], promoting open access and enhancing reproducibility.